# Regular functional covering numbers

### v1

@@ 3 Regular functional covering numbers
$N(\mathds{1}_{a(k+1)r_{f}B_{2}^{n}},\mathds{1}_{R_{f}})\leqslant\bigl(2\sqrt{t}\,a(k+1)+1\bigr)^{n}\exp\!\left(\frac{c\delta_{n}^{2}n}{t}\right),$
@@
$N(g,\mathds{1}_{R_{f}})\leqslant\exp\!\left(\frac{c\delta_{n}^{2}n}{t}\right)\sum_{k=0}^{\infty}\bigl(e^{-c_{0}a^{2}k^{2}+2\sqrt{t}\,a(k+1)}\bigr)^{n}.$
Choosing $a=c_{2}\sqrt{t}$ (for an absolute constant $c_{2}>0$ sufficiently large) makes the series bounded by an absolute constant. Hence
$N(g,\mathds{1}_{R_{f}})\leqslant C\exp\!\left(\frac{c\delta_{n}^{2}n}{t}\right).$
@@
Write $f=e^{-\varphi}$, where $\varphi$ is convex. Then $f^{\ast}=e^{-\mathcal{L}\varphi}$, where $\mathcal{L}\varphi$ is the Legendre transform. By [15, Lemma 8], for all $s,t>0$,
(3.22) $t\{x:\varphi(x)\leqslant t\}^{\circ}\;\subset\;\{y:\mathcal{L}\varphi(y)\leqslant t\}\;\subset\;(t+s)\{x:\varphi(x)\leqslant s\}^{\circ}.$
@@
$N(R_{f^{\ast}},\,tr^{\ast}B_{2}^{n})\leqslant N(100n\,(R_{f})^{\circ},\,50nt\,r^{\circ}B_{2}^{n})\leqslant\exp\!\left(\frac{c\gamma_{n}^{2}n}{t^{2}}\right),$
@@
From Lemma 3.3 we know that $R_{f}$ has bounded geometric distance from a centered convex body. Therefore, we may apply the Blaschke-Santaló and Bourgain-Milman inequalities (up to an absolute constant) to $R_{f}$. Combining with (3.23) yields
$\mathrm{vol}_{n}(R_{f^{\ast}})^{1/n}\approx n\,\mathrm{vol}_{n}((R_{f})^{\circ})^{1/n}\approx 1\Big/\mathrm{vol}_{n}(R_{f})^{-1/n}\approx 1,$
@@
The convex body $K_{n+1}(f)$ is isotropic, and therefore, by Proposition 2.1,
@@
Combining (3.6) with (3.24) and (3.25), we deduce that $R_{f}$ satisfies (3.8) and (3.16). Moreover,
$\mathrm{vol}_{n}(R_{f})^{1/n}\approx\mathrm{vol}_{n}(K_{n+1}(f))^{1/n}=1.$
Applying Propositions 3.4 and 3.5, we obtain
$N(f,\,t\odot g)\leqslant\exp\!\left(\frac{c\gamma_{n}^{2}n}{t}\right)\quad\text{and}\ N(g,\,t\odot f)\leqslant\exp\!\left(\frac{c\delta_{n}^{2}n}{t}\right)\qquad\text{for all }t\geqslant 1.$
@@
Applying Propositions 3.4 and 3.5 to $f^{\ast}$ yields
$N(f^{\ast},\,t\odot g)\leqslant\exp\!\left(\frac{c\delta_{n}^{2}n}{t}\right)\quad\text{and}\quad N(g,\,t\odot f^{\ast})\leqslant\exp\!\left(\frac{c\gamma_{n}^{2}n}{t}\right)\qquad\text{for all }t\geqslant 1.$
@@
For the proof of Theorem 1.5 we shall use the next lemma from [24].
@@
$\mathrm{vol}_{n}(R_{f})^{1/n}\approx\mathrm{vol}_{n}(K_{n+1}(f))^{1/n}=1$
and $R_{f}$ satisfies (3.8) and (3.16).
Combining (3.28) with the Blaschke-Santaló and Bourgain-Milman inequalities we get
@@
$N(R_{f_{\mathcal{A}}},\,tr_{\mathcal{A}}B_{2}^{n})\leqslant N(2n\,(R_{f})^{\circ},\,tn\,r^{\circ}B_{2}^{n})\leqslant\exp\!\left(\frac{c\gamma_{n}^{2}n}{t^{2}}\right),$
@@
Since $R_{f_{\mathcal{A}}}$ satisfies (3.20), and moreover
@@
applying Propositions 3.4 and 3.5 to $f_{\mathcal{A}}$ yields
$N(f_{\mathcal{A}},\,t\odot g)\leqslant\exp\!\left(\frac{c\delta_{n}^{2}n}{t}\right)\quad\text{and}\quad N(g,\,t\odot f_{\mathcal{A}})\leqslant\exp\!\left(\frac{c\gamma_{n}^{2}n}{t}\right)\qquad\text{for all }t\geqslant 1.$
@@
We begin as in the proof of Theorem 1.3. Recall that $K_{n+1}(f)$ is isotropic, and hence satisfies (3.29). Since $R_{f}$ is at bounded geometric distance from $K_{n+1}(f)$, we obtain $r_{f}\approx\sqrt{n}$, or equivalently,
$\mathrm{vol}_{n}(R_{f})^{1/n}\approx\mathrm{vol}_{n}(K_{n+1}(f))^{1/n}=1,$
@@
$N(r^{\ast}B_{2}^{n},\,R_{f^{\ast}})\leqslant N(c_{2}n\,r^{\circ}B_{2}^{n},\,c_{1}n\,(R_{f})^{\circ})\leqslant C_{4}^{n}.$
@@
Acknowledgement. The second named author acknowledges support by a PhD scholarship from the National Technical University of Athens.

### v2

@@ 3 Regular functional covering numbers
$N(\mathds{1}_{a(k+1)r_{f}B_{2}^{n}},\mathds{1}_{R_{f}})\leqslant\bigl(2\sqrt{\kappa_{n}t}\,a(k+1)+1\bigr)^{n}\exp\!\left(\frac{c\kappa_{n}n}{t}\right),$
@@
$N(g,\mathds{1}_{R_{f}})\leqslant\exp\!\left(\frac{c\kappa_{n}n}{t}\right)\sum_{k=0}^{\infty}\bigl(e^{-c_{0}a^{2}k^{2}+2\sqrt{\kappa_{n}t}\,a(k+1)}\bigr)^{n}.$
Choosing $a=c_{2}\sqrt{\kappa_{n}t}$ (for an absolute constant $c_{2}>0$ sufficiently large) makes the series bounded by an absolute constant. Hence
$N(g,\mathds{1}_{R_{f}})\leqslant C\exp\!\left(\frac{c\kappa_{n}n}{t}\right).$
@@
Remark 3.8.
(i) In the proof of Proposition 3.7 we used only the bound $N(r_{f}B_{2}^{n},tR_{f})\leqslant\exp\left(\tfrac{c_{1}\kappa_{n}^{2}n}{t^{2}}\right)$ from (3.18). However, note that in (3.19) we need to assume that $f(0)=\|f\|_{\infty}=1$, i.e. that $f$ is a geometric log-concave function.
(ii) Instead of decomposing $g(x)$ into a sum, as in (3.21), one may write
$g(x)=\exp\left(-\tfrac{1}{2}|x|^{2}\right)=\int_{0}^{\infty}se^{-s^{2}/2}\mathds{1}_{sB_{2}^{n}}(x,s)\,ds$
and then estimate the functional covering number $N(g,\mathds{1}_{R_{f}})$ writing
$N(g,\mathds{1}_{R_{f}})\leqslant\int_{0}^{\infty}se^{-s^{2}/2}N(\mathds{1}_{sB_{2}^{n}},\mathds{1}_{R_{f}})\,ds.$
Making the change of variables $s=aur_{f}$ and using the same argument as in the proof above, we get
$N(\mathds{1}_{aur_{f}B_{2}^{n}},\mathds{1}_{R_{f}})\leqslant\bigl(2\sqrt{\kappa_{n}t}\,au+1\bigr)^{n}\exp\!\left(\frac{c\kappa_{n}n}{t}\right),$
and recalling that $r_{f}\approx\sqrt{n}$ we see that
$N(g,\mathds{1}_{R_{f}})\leqslant\exp\!\left(\frac{c\kappa_{n}n}{t}\right)\int_{0}^{\infty}\bigl(e^{-c_{0}a^{2}u^{2}+3\sqrt{\kappa_{n}t}\,au}\bigr)^{n}du.$
Choosing $a=c_{2}\sqrt{\kappa_{n}t}$ (for an absolute constant $c_{2}>0$ sufficiently large) makes the last integral bounded by an absolute constant, and Proposition 3.7 follows.
@@
Write $f=e^{-\varphi}$, where $\varphi$ is a convex geometric function. Then $f^{\ast}=e^{-\mathcal{L}\varphi}$, where $\mathcal{L}\varphi$ is the Legendre transform of $\varphi$. By [15, Lemma 8], for all $s,t>0$,
(3.26) $t\{x:\varphi(x)\leqslant t\}^{\circ}\;\subseteq\;\{y:\mathcal{L}\varphi(y)\leqslant t\}\;\subseteq\;(t+s)\{x:\varphi(x)\leqslant s\}^{\circ}.$
@@
$N(R_{f^{\ast}},\,tr^{\ast}B_{2}^{n})\leqslant N(100n\,(R_{f})^{\circ},\,50nt\,r^{\circ}B_{2}^{n})\leqslant\exp\!\left(\frac{c\delta_{n}^{2}n}{t^{2}}\right),$
@@
From Lemma 3.3 we know that $R_{f}$ has bounded geometric distance from a centered convex body. Therefore, we may apply the Blaschke–Santaló and Bourgain–Milman inequalities (up to an absolute constant) to $R_{f}$. Combining with (3.27) yields
$\mathrm{vol}_{n}(R_{f^{\ast}})^{1/n}\approx n\,\mathrm{vol}_{n}((R_{f})^{\circ})^{1/n}\approx\mathrm{vol}_{n}(R_{f})^{-1/n}\approx 1,$
@@
Remark 3.10.
The precise statement of [15, Lemma 8] is that if $\varphi:\mathbb{R}^{n}\to(-\infty,\infty]$ is a convex function then, for all $s,t>0$,
(3.28) $t\{x:\varphi(x)\leqslant t+\varphi(0)\}^{\circ}\;\subseteq\;\{y:\mathcal{L}\varphi(y)\leqslant t-\min\varphi\}\quad\text{and}\quad\{y:\mathcal{L}\varphi(y)\leqslant t\}\;\subseteq\;(t+s)\{x:\varphi(x)\leqslant s\}^{\circ}.$
Assuming that $f$ is a centered log-concave function with $\|f\|_{\infty}=1$, we have $\min\varphi=0$ and $e^{-\varphi(0)}=f(0)\geqslant e^{-n}\|f\|_{\infty}=e^{-n}$ (by Fradelizi’s inequality) which implies $\varphi(0)\leqslant n$. Setting $s=t=50n$ in (3.28) we get
$50n(\tilde{R}_{f})^{\circ}\subseteq R_{f^{\ast}}\subseteq 100n(R_{f})^{\circ},$
where $\tilde{R}_{f}=R_{51n}(f)$. Log-concavity of $f$ implies that $\tilde{R}_{f}\subseteq\frac{51}{50}R_{f}$, and hence
$R_{f^{\ast}}\approx n(R_{f})^{\circ}.$
Then, we can follow the above argument and conclude that Proposition 3.9 holds true for every centered log-concave function $f:\mathbb{R}^{n}\to[0,\infty)$ for which $\|f\|_{\infty}=1$, $\mathrm{vol}_{n}(R_{f})^{1/n}\approx 1$ and $R_{f}$ satisfies (3.24) (the assumption that $f(0)=\|f\|_{\infty}$ is not needed).
@@
The convex body $K_{n+1}(f)$ is almost isotropic, and Remark 3.4 shows that
@@
Combining (3.7) with (3.29) and (3.30), we deduce that $R_{f}$ satisfies (3.9) with $\kappa_{n}=\gamma_{n}$ and (3.18) with $\kappa_{n}=\delta_{n}$. Moreover,
$\mathrm{vol}_{n}(R_{f})^{1/n}\approx\mathrm{vol}_{n}(K_{n+1}(f))^{1/n}\approx 1.$
Applying Propositions 3.5 and 3.7, we obtain
$N(f,\,t\odot g)\leqslant\exp\!\left(\frac{c\gamma_{n}n}{t}\right)\quad\text{and}\ N(g,\,t\odot f)\leqslant\exp\!\left(\frac{c\delta_{n}n}{t}\right)\qquad\text{for all }t\geqslant 1.$
@@
Applying Propositions 3.5 and 3.7 to $f^{\ast}$ yields
$N(f^{\ast},\,t\odot g)\leqslant\exp\!\left(\frac{c\delta_{n}n}{t}\right)\quad\text{and}\quad N(g,\,t\odot f^{\ast})\leqslant\exp\!\left(\frac{c\gamma_{n}n}{t}\right)\qquad\text{for all }t\geqslant 1.$
@@
For the proof of Theorem 1.5 we shall use the next result of V. Milman and Rotem (see [24, Proposition 11]).
@@
$\mathrm{vol}_{n}(R_{f})^{1/n}\approx\mathrm{vol}_{n}(K_{n+1}(f))^{1/n}\approx 1$
and $R_{f}$ satisfies (3.9) with $\kappa_{n}=\gamma_{n}$ and (3.18) with $\kappa_{n}=\delta_{n}$.
Combining (3.33) with the Blaschke–Santaló and Bourgain–Milman inequalities we get
@@
$N(R_{f_{\mathcal{A}}},\,tr_{\mathcal{A}}B_{2}^{n})\leqslant N(2n\,(R_{f})^{\circ},\,tn\,r^{\circ}B_{2}^{n})\leqslant\exp\!\left(\frac{c\delta_{n}^{2}n}{t^{2}}\right),$
@@
Since $R_{f_{\mathcal{A}}}$ satisfies these bounds, and moreover
@@
applying Proposition 3.5 with $\kappa_{n}=\delta_{n}$ and Proposition 3.7 with $\kappa_{n}=\gamma_{n}$ to $f_{\mathcal{A}}$ yields
$N(f_{\mathcal{A}},\,t\odot g)\leqslant\exp\!\left(\frac{c\delta_{n}n}{t}\right)\quad\text{and}\quad N(g,\,t\odot f_{\mathcal{A}})\leqslant\exp\!\left(\frac{c\gamma_{n}n}{t}\right)\qquad\text{for all }t\geqslant 1.$
@@
We begin as in the proof of Theorem 1.3. Recall that $K_{n+1}(f)$ is almost isotropic (see Remark 3.4), and hence satisfies (3.34). Since $R_{f}$ is at bounded geometric distance from $K_{n+1}(f)$, we obtain $r_{f}\approx\sqrt{n}$, or equivalently,
$\mathrm{vol}_{n}(R_{f})^{1/n}\approx\mathrm{vol}_{n}(K_{n+1}(f))^{1/n}\approx 1,$
@@
$N(r^{\ast}B_{2}^{n},\,R_{f^{\ast}})\leqslant N(c_{3}n\,r^{\circ}B_{2}^{n},\,c_{4}n\,(R_{f})^{\circ})\leqslant C_{4}^{n}.$
@@
Acknowledgement. We are grateful to the referee for a careful and thoughtful reading of the manuscript. Their insightful comments and constructive suggestions have helped us weaken the assumptions and improve the estimates in some of the results. The second named author acknowledges support by a PhD scholarship from the National Technical University of Athens.